OpenStudy (anonymous):

Sam bought two CDs for $13 each.Sales tax for both CDs was $3. Write an expression to show how much Sam paid in all.

OpenStudy (anonymous):

ok ill help

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

13*2= 26 + 6 =32 $ work with that

OpenStudy (anonymous):

or you can do this one

OpenStudy (anonymous):

13*X+6=32 or something like that

OpenStudy (anonymous):

More or less I'm kinda confused here but no offence through

OpenStudy (anonymous):

Xd i cant remember how to write expression... i feel dumb

OpenStudy (anonymous):

y= 13x + 3 Sorry skler but I have no idea what you're doing

OpenStudy (anonymous):

THATS WHAT I WAS GOING FOR

OpenStudy (anonymous):

thanks guy

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

Y = total cost X = number of cds bought

OpenStudy (anonymous):

yea forgot*facepalm*

OpenStudy (anonymous):

So what are am I supposed to do again?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

XD y= 13x + 3

OpenStudy (anonymous):

is what you need to put

OpenStudy (anonymous):

y = 13x + 3

OpenStudy (anonymous):

So u mean put 13 times +3?